# TaskEval: Assessing Difficulty of Code Generation Tasks for Large Language Models

### v2

@@ 5.1. RQ1: How do tasks’ characteristics varies across benchmark?
Figure 3. Maps of the difficulty $\delta_{j}$ vs discriminant $a_{j}$ of each task for a give benchmark. The color represent the expected probability obtained on a given task by a hypothetical CodeLLM which capacity $\overline{\theta}$ is the average of the capacity of our CodeLLMs: (Left) HumanEval+ tasks, (Right) ClassEval tasks.
[…]
@@
Figure 4. Proportion of tasks in HumanEval+ (Top) and ClassEval (Bottom) with a given expected probability of response for each CodeLLM.
@@
We now present the mapping of difficulty and discriminant parameters for each task in both benchmarks, as computed by the IRT models, in Figure 3. For each task, we also report the expected probability $E[p_{ij}|\overline{\theta},\delta_{j},a_{j}]$ as predicted by the model for a hypothetical CodeLLM with a capacity $\overline{\theta}$ as the average of all our CodeLLMs. For both benchmark, CodeLLMs obtained capacity $\theta_{i}$ follows the outcome we could expect: GPT3.5 is the most capable CodeLLM, closely followed by Deepseek and MagiCoder that are tied together, then Code Gemma and finally Code Llama. From the mapping, we first observe a similar general shape on the positive discriminant side with some forms of tail on the negative discriminant side. Secondly, still on the positive discriminant side, we see that the difficulty correlates with the expected probability: intuitively, the higher the difficulty of the task, the smaller the probability of answer should be. We also observe some “peak” of high discriminability around tasks of difficulty between $0.2$ and $0.4$ for both benchmarks though the trend is more pronounced for HumanEval+: the most discriminative tasks, in the sense allowing to tell which CodeLLMs are better, are more likely to be the middle of the range of difficulty. Indeed, for hard tasks or easy tasks, either all CodeLLMs fail or all CodeLLMs manage in a similar proportion, which makes it difficult to discriminate between them. This echoes what we observed with the cumulative distribution of the expected probability of answer in Figure 4 with a wider discrepancy between CodeLLMs.
@@
We note however an inverse pattern in terms of difficulty / expected probability when we reach the tasks with a negative discriminant. It should be noted, from Figure 2, that a negative discriminant means that less capable CodeLLMs are more likely to have a correct answer than more capable CodeLLMs. We make the distinctions between two parts of this side of the diagram: the tasks with high difficulty and low difficulty as symbolized by the $0.5$ threshold. For tasks with negative discriminant and high difficulty, we also note that CodeLLMs tend to also have a relatively high expected probability on these tasks. In that case, this can be explained by both non-determinism in the prompts and the small number of CodeLLMs used. For instance, in the green curve in Figure 2, if all CodeLLMs have abilities between $0.3$ and $0.5$, the difference between expected probability would be small but the IRT models would yield a negative discriminant with a high difficulty. In that case, it is even possible that, by adding more CodeLLMs, that the slope could be rectified or even inverted. The second part of this half of the diagrams have tasks with low difficulty and negative discriminant. In that case, most CodeLLMs will fail such tasks. This is generally symptomatic of possible annotation errors as less capable CodeLLM do better than more capable ones but all CodeLLMs are generally not good. We thus manually investigated these tasks. Among the $9$ tasks from ClassEval and $3$ tasks from HumanEval+, we found issues in the way the original prompts (and so our prompts generated from them) were formulated for $7$ of them in ClassEval and $1$ in HumanEval+. For instance, in ClassEval, one task consists in implementing the cosine function using a taylor approximation. We give in Listing 4 an excerpt with relevant information. From the excerpt, the CodeLLMs have access to the signature of the function implementing the ‘taylor‘ function but not the actual implementation. The oracle implementation contains codes to convert $x$ into radians before returning the approximation. This poses an issue as without this information, most CodeLLMs will generate codes in the ‘cos‘ that convert $x$ into radian before using the ‘taylor‘ function, leading to an error. Code LLama and Code Gemma tend not to follow this approach as much, hence why the negative discriminant and low expected probability.
@@ 14 Calculate the cos value of the x-degree angle
17 $>$ tricalculator = TriCalculator()
18 $>$ tricalculator.cos(60)
@@ 17 $>$ tricalculator = TriCalculator()
Listing 4: Excerpt from the ‘cos‘ task of ClassEval following the original prompt
On the other hand, for the tasks we investigated where no issues were detected, we observed behaviors that may reveal specific characteristics of CodeLLMs, possibly influenced by their training data. Given in Listing 5 is the original prompt for a task of HumanEval+ where the goal is to filter a list to retain only integers. While relatively simple, more able CodeLLMs like GPT3.5 fail several times on it when Code LLama is better at managing it. Looking at the answers provided, the codes generated are similar, except that GPT3.5 generally relies on ‘isinstance‘ while Code LLama relies on ‘type‘ to check for integers. While ‘isinstance‘ is generally more useful as it accounts for inheritance, however, in that case, ‘type‘ should be used. Indeed, if the list would, for instance, contain boolean ‘True‘ or ‘False‘, ‘isinstance‘ would consider them as integer (as boolean type inherits from integer in Python) while ‘type‘ will not. This can be indicative of specificity of the training dataset or, in the worst case, some form of memorization.
@@ 5 Filter given list of any python values only for integers
Listing 5: Excerpt from the ‘filter_integers‘ task of HumanEval+ following the original prompt
Findings 1: IRT allows us to map tasks in terms of difficulty and discriminant and to contrast benchmarks. CodeLLMs have varying behavior within and between benchmarks: ClassEval tends to have more extreme (very easy/difficult) tasks while HumanEval+ have more medium tasks in terms of difficulty for the CodeLLMs under study, with more discrepancies in terms of expected answers in HumanEval+. Most discriminant tasks in both benchmarks generally have a difficulty between $0.2$ and $0.4$, which thus are better at separating CodeLLMs. Few tasks have negative discriminant, symptomatic of small differences due to non-determinism or annotation errors.
@@ 5.2. RQ2: Do tasks’ topics affect their characteristics?
In this section, we proceed with the exploration of the topics. Results are given in Figure 5. Starting with HumanEval+, we see that most topics have a mean difficulty in the easy-medium range ($0.3-0.5$). Only the last three topics show a high variation from the rest of the topics, however it might be explained by the lower number of individual tasks in these topics making them particularly specialized. Out of the other topics, three of them have a mean difficulty above $0.45$: Nested parentheses/brackets (12), Sequences (8), and Strings to number (5). Sequences (e.g., Fibonacci) also have a high positive discriminant, pointing out that less capable CodeLLMs such as Code Llama tend to have a hard time processing those tasks. Such tasks can be interesting to distinguish between different CodeLLMs. On the contrary, Strings to number (e.g., converting roman numeral to numbers) has a mean discriminant lower than 1, showing that CodeLLMs perform somewhat similarly on such tasks. The mean accuracy of the CodeLLMs varies from topic to topic, but $GPT3.5$ is generally the model with the highest mean accuracy and Code Llama the one with the lowest. Most topics have a mean discriminant around $1$ and no topics have a negative mean discriminant. Such topics are more likely to distinguish between more/less capable CodeLLMs, though there can be some variability between similarly capable CodeLLMs such as MagiCoder and Deepseek or Code Llama and Code Gemma.
@@
Figure 5. Mean per topic of: 1) the accuracy on the tasks’ prompts for each CodeLLM, 2) the tasks’ difficulty (Diff), 3) the tasks’ discriminant (Disc). (Top) HumanEval+, (Bottom) ClassEval.
@@
Regarding ClassEval, we see that contrary to HumanEval+ most topics are fairly easy with $48\%$ of the topics with a mean difficulty below $0.3$ and most CodeLLMs achieving a higher mean accuracy compared to HumanEval+. As such, Statistics (7) (e.g., calculating a mean), Files reading (13) (e.g., opening a JSON file and returning the content), and Data Structure information retrieval (17) (e.g., finding data in a dictionary) are relatively trivial for all CodeLLMs. On the contrary, besides Temperature processing (18), which might have a very high difficulty due to a low number of very specialized tasks (dealing with classes with functions related to ambient temperature control), four topics have a mean difficulty above $0.45$: Data structure calculations (3) (e.g., computing Pearson correlations between two lists), SQL requests (8) (see the example in Listing 1, Section 2), Text encryption/decryption (14) (e.g., implementing Caesar cipher), and Strings replacements (16) (e.g., replacing a character by another in a string). Compared to HumanEval+, there are more low discriminant topics (discriminant $<1$ or even $<0.75$) which highlights that ClassEval contains more tasks where CodeLLMs tend to obtain similar results. As we mentioned in Section 5.1, there were fewer discrepancies between CodeLLMs in terms of the expected probability of answer in ClassEval than HumanEval+. Within these topics, we can even see some of them for which less capable CodeLLMs (Code Gemma and Code Llama) achieve a higher mean accuracy than more capable ones: it is the case for String processing (4) (e.g., removing stop-words from a text with NLTK), Data structure items additions (15), and to a lesser degree SQL requests (8). The case of Data structure items additions (15) is interesting as it is the only topic for which Code Llama and Code Gemma are on average better than even GPT3.5, as highlighted by the average negative discriminant. Analyzing the tasks within the topic, we observe that more capable models often infer additional assumptions about the task, leading to unnecessarily complex generated code. As an example, we give the generated codes for one such task of GPT3.5 and Code Llama along with the original prompt in Listing 6.
@@
Listing 6: Example of discrepancy on a task from Topic 15. We give first one of our prompt for the task, then generated code by GPT3.5, and CodeLlama.
As one can see, GPT3.5 assumes there is a need to check for some conflicts, possibly due to the ‘course‘ parameter containing information such as ‘start_time‘ and the function ‘self.check_course_conflict‘ being available as a class function from the prompt. However, this is not requested at all in the prompt (and in the task), which is quite straightforward. On the contrary, Code Llama manages the task by following exactly the instructions mentioned in the prompt. We observe a similar phenomenon for other generated code on the same task. This behaviour explains the negative discriminant on the task: less capable CodeLLMs such as Code LLama end up having better performance than more able CodeLLMs which tend to overdo it when addressing the task.
Findings 2: Analyzing topics can reveal variations in how CodeLLMs handle different types of tasks. Some topics are notably harder for CodeLLMs in both benchmarks (e.g., Sequences and Nested parentheses/brackets in HumanEval+ and SQL requests in ClassEval) while some are trivial (e.g., Statistics in ClassEval). Discriminant information gives us clues on some topics of interest, highlighting tasks where CodeLLMs behave singularly.

### v3

@@ 5. Results
5.1. Validation
Before analysing our results, we first proceed to validate the generated prompts. To validate our generated prompts, on top of a manual check, we first inspect the semantic similarity of the prompts rephrasing within and across levels. To do so, we used BERTScore (Zhang et al., 2019), a semantic similarity metric leveraging the cosine similarity of the embedding of references/candidate solutions through a transformer. We did so for each rephrasing against the original prompt of the same level and of other levels. The obtained median similarity for HumanEval+ ranges from 0.60 (rephrasing of Level 3) to 0.71 (rephrasing of Level 1), and for ClassEval ranges from 0.60 (rephrasing of Level 2 and Level 3) to 0.65 (rephrasing of Level 1). For some rephrasing, similarity can drop as low as 0.38 (HumanEval+) or 0.28 (ClassEval). Inspecting these cases, we did not observe semantic discrepancies. However, the rephrasing obtained via GPT can sometimes be quite convoluted, which might artificially push semantic similarity down, on top of BERTScore’s inherent limitation (Hanna and Bojar, 2021). For instance, “Clears the expression of all characters that are not brackets.” and “Retain only bracket characters in the expression, removing all others.” leads to a similarity score of 0.28 despite both sentences expressing the same idea. To further evaluate the validity of the samples, we compared the average similarity within and across levels between rephrasing for the same task using the Wilcoxon test (Wilcoxon, 1992), a nonparametric statistical test evaluating the null hypothesis that two paired samples come from the same distribution. In all cases, we obtained a statistically significant difference (p-value ¡ 0.01) with a strong effect size (¿ 0.8) (Fritz et al., 2012) when testing whether the distribution of same-level rephrasings was greater than the across-level rephrasings. This highlights that same-level rephrasings are semantically closer to each other than they are to other rephrasings of other levels for the same task.
To further verify the validity of the prompts, we retrained independent IRT models for both datasets using only the aggregated score of code samples of a single level, resulting in 6 different IRT models (3 for each dataset). We would expect that the difficulty of the tasks will drop as the prompt level of information increases. For HumanEval+, we found that the median difficulty of the tasks went from 0.42 (Level 1) to 0.15 (Level 3). Similarly, for ClassEval, the median difficulty of the tasks dropped from 0.27 (Level 1) to 0.05 (Level 3). Further, using the statistical test of Wilcoxon showed a statistically significant difference (p-value ¡ 0.01) with effect size ranging from 0.25 to 0.64 (Fritz et al., 2012), when comparing the difficulty for the same task across IRT models of different levels. The lowest effect size was obtained when comparing the difficulty of tasks between Level 1 and Level 2, and the highest was obtained when comparing Level 1 and Level 3. This validates that a higher level of information results in lower difficulty of the tasks, as LLMs get increasingly more information, which allows for more correct samples to be generated.
@@ 5.2. RQ1: How do tasks’ characteristics varies across benchmark?
Figure 3. Proportion of tasks in HumanEval+ (Top) and ClassEval (Bottom) with a given expected probability of response for each CodeLLM.: Cumulative distribution of HumanEval+ and ClassEval frequency of tasks with respect to expected answers differs
We first calculated the cumulative distribution of the expected probability of having a correct answer, based on IRT modelling, for each task in each benchmark. The results are presented in Figure 3 and can be read as follows: for instance, for HumanEval+, $30\%$ of the tasks for DeepSeek-V3 have an expected probability of a correct answer of $90\%$ or less (and so $70\%$ have probability $90\%$ or more). These graphs allow us to compare how, globally, tasks are handled by CodeLLMs in both benchmarks. The first observation we can make is regarding the discrepancies between CodeLLMs inside the same benchmark. In the case of HumanEval+, there are wider variations between CodeLLMs compared to ClassEval. Notably, Code Gemma, Code LLama, and Yi-Coder seem to struggle more on HumanEval+ with less than $\sim 50\%$ chance of generating a correct code on over $40\%$ of the tasks, compared to the other CodeLLMs. In a second time, we can observe the difference for the same CodeLLM across benchmarks. We observe that ClassEval tends to have more extreme tasks compared to HumanEval+. Indeed, there are more very easy ($90\%$ or more probability of correct answer) and very hard tasks ($10\%$ or less probability of correct answer) for ClassEval compared to HumanEval+ for most CodeLLMs. On the contrary, HumanEval+ have more, on average, medium tasks (probability of correct answer between $40\%$ and $60\%$). The difference tends to narrow for models extreme, mainly CodeLLama and DeepSeek-V3, on ClassEval compared to HumanEval+. Overall, for the same CodeLLM, there is a difference among benchmarks in terms of the expected probability of answer: computing the AD test returns a statistically significant difference (pvalue<0.05-value<0.05) in all cases.
Figure 4. Maps of the difficulty $\delta_{j}$ vs discriminant $a_{j}$ of each task for a give benchmark. The colour represents the expected probability obtained on a given task by a hypothetical CodeLLM whose capacity $\overline{\theta}$ is the average of the capacity of our CodeLLMs: (Left) HumanEval+ tasks, (Right) ClassEval tasks.: HumanEval+ and ClassEval exhibit similar patterns when plotting difficulty against discriminant of tasks
@@
We now present the mapping of difficulty and discriminant parameters for each task in both benchmarks, as computed by the IRT models, in Figure 4. For each task, we also report the expected probability $E[p_{ij}|\overline{\theta},\delta_{j},a_{j}]$ as predicted by the model for a hypothetical CodeLLM with a capacity $\overline{\theta}$ as the average of all our CodeLLMs. For both benchmarks, CodeLLMs obtained capacity $\theta_{i}$ follow the outcome we could expect: DeepSeek-V3 is the most capable CodeLLM, followed by GPT3.5, then DeepSeekCoder, MagiCoder, and QwenCoder are tied together, and finally Code Gemma, Code Llama, and Yi-Coder are similar. From the mapping, on the positive discriminant side, we see that the difficulty correlates with the expected probability: intuitively, the higher the difficulty of the task, the smaller the probability of the answer should be. We also observe some “peak” of high discriminability around tasks of difficulty between $0.4$ and $0.6$ for HumanEval+ and around $0.2$ and $0.4$ as well as $0.6$ and $0.8$ for ClassEval: the most discriminative tasks, allowing to tell apart which CodeLLMs are better, are more likely to not to be in the extreme of difficulty. Indeed, for very hard tasks or very easy tasks, either all CodeLLMs fail or all CodeLLMs manage in a similar proportion, which makes it difficult to discriminate between them. This echoes what we observed with the cumulative distribution of the expected probability of answer in Figure 3, with a wider discrepancy between CodeLLMs.
We note, however, an inverse pattern in terms of difficulty / expected probability when we reach the tasks with a negative discriminant. It should be noted, from Figure 2, that a negative discriminant means that less capable CodeLLMs are more likely to have a correct answer than more capable CodeLLMs. We make the distinctions between two parts of this side of the diagram: the tasks with high difficulty and low difficulty, as symbolised by the $0.5$ threshold. For tasks with negative discriminant and high difficulty, we also note that CodeLLMs tend to have a relatively high expected probability on these tasks. In that case, this can be explained by both non-determinism in the prompts and the small number of CodeLLMs used. For instance, in the green curve in Figure 2, if all CodeLLMs have abilities between $0.3$ and $0.5$, the difference between expected probabilities would be small, but the IRT models would yield a negative discriminant with a high difficulty. We noted this behaviour when realising our experiments: computing the difficulty/discriminant with a subset of models compared to the whole set resulted in some increase of the number of tasks with a negative discriminant, with the magnitude of the increase varying depending on the subset chosen. For instance, when using models {deepseekcoder, magicoder, codellama, codegemma, gpt} only, we observed an increase from 1.2% to 8.5% on HumanEval+ and 9.5% to 15% on ClassEval of the proportions of such tasks compared to using all models.
@@
The second part of this half of the diagrams has tasks with low difficulty and a negative discriminant. In that case, most CodeLLMs will fail such tasks. This is generally symptomatic of possible annotation errors, as less capable CodeLLMs do better than more capable ones, but not all CodeLLMs are generally good. We thus manually investigated these tasks. Among the $9$ tasks from ClassEval and $1$ task from HumanEval+, we found issues in the way the original prompts (and so our prompts generated from them) were formulated for $7$ of them in ClassEval and $1$ in HumanEval+. For instance, in ClassEval, one task consists of implementing the cosine function using a Taylor approximation. We give in Listing 5 an excerpt with relevant information. From the excerpt, the CodeLLMs have access to the signature of the function implementing the ‘taylor‘ function, but not the actual implementation. The oracle implementation contains code to convert $x$ into radians before returning the approximation. This poses an issue as, without this information, most CodeLLMs will generate codes in the ‘cos‘ that convert $x$ into radians before using the ‘taylor‘ function, leading to an error. Code LLama and Code Gemma tend not to follow this approach as much, hence why the a negative discriminant and a low expected probability.
@@ 14 Calculate the cos value of the x-degree angle
17 >>> tricalculator = TriCalculator()
18 >>> tricalculator.cos(60)
@@
Listing 5: Excerpt from the ‘cos‘ task of ClassEval following the original prompt
On the other hand, for the tasks we investigated where no issues were detected, we observed behaviours that may reveal specific characteristics of CodeLLMs, possibly influenced by their training data. Given in Listing 6 is the original prompt for a task of HumanEval+ where the goal is to filter a list to retain only integers. While relatively simple, more able CodeLLMs like GPT3.5 fail several times on it when Code LLama is better at managing it. Looking at the answers provided, the codes generated are similar, except that GPT3.5 generally relies on ‘isinstance‘ while Code LLama relies on ‘type‘ to check for integers. While ‘isinstance‘ is generally more useful as it accounts for inheritance, however, in that case, ‘type‘ should be used. Indeed, if the list were, for instance, to contain boolean ‘True‘ or ‘False‘, ‘isinstance‘ would consider them as integers (as the boolean type inherits from integer in Python) while ‘type‘ will not. This can be indicative of the specificity of the training dataset or, in the worst case, some form of memorisation.
@@ 5 Filter given list of any python values only for integers
Listing 6: Excerpt from the ‘filter_integers‘ task of HumanEval+ following the original prompt
Findings 1: IRT allows us to map tasks in terms of difficulty and discriminant and to contrast benchmarks. CodeLLMs have varying behaviour within and between benchmarks: ClassEval tends to have more extreme (very easy/difficult) tasks while HumanEval+ have more medium tasks in terms of difficulty for the CodeLLMs under study, with more discrepancies in terms of expected answers in HumanEval+. Most discriminant tasks in both benchmarks generally have a difficulty between $0.2$ and $0.8$, which thus are better at separating CodeLLMs. Few tasks have negative discriminant, symptomatic of small differences due to non-determinism or annotation errors.
@@ 5.3. RQ2: Do tasks’ topics affect their characteristics?
In this section, we proceed with the exploration of the topics. Results are given in Figure 5. Starting with HumanEval+, we see that most topics have a mean difficulty in the easy-medium range (0.30.50.3-0.5). The last two topics show a high variation from the rest of the topics; however, this might be explained by the lower number of individual tasks in these topics, making them particularly specialised. Out of the other topics, four of them have a mean difficulty above $0.45$: Nested parentheses/brackets (12), Sequences (8), Special patterns in strings (4) and Replacements in strings (14). Sequences (e.g., Fibonacci) also have a high positive discriminant, pointing out that less capable CodeLLMs such as Code Llama tend to have a hard time processing those tasks. Such tasks can be interesting in distinguishing between different CodeLLMs. The mean accuracy of the CodeLLMs varies from topic to topic, but DeepSeek-V3 is generally the model with the highest mean accuracy and Code Llama the one with the lowest. Most topics have a mean discriminant around $1$, and no topics have a negative mean discriminant. Such topics are more likely to distinguish between more/less capable CodeLLMs, though there can be some variability between similarly capable CodeLLMs such as MagiCoder, Deepseek and QwenCoder or Code Llama, Code Gemma and YiCoder.
@@
Figure 5. Mean per topic of: 1) the accuracy on the tasks’ prompts for each CodeLLM, 2) the tasks’ difficulty (Diff), 3) the tasks’ discriminant (Disc). (Top) HumanEval+, (Bottom) ClassEval.: Difficulty/Discriminant of extracted topics varies for both ClassEval and HumanEval+
@@
Regarding ClassEval, we see that, contrary to HumanEval+, most topics are fairly easy, with $48\%$ of the topics with a mean difficulty below $0.3$ and most CodeLLMs achieving a higher mean accuracy compared to HumanEval+. As such, SQL Requests (8), Files reading (13) (e.g., opening a JSON file and returning the content), and Data Structure information retrieval (17) (e.g., finding data in a dictionary) are relatively trivial for all CodeLLMs. On the contrary, besides Temperature processing (18), which might have a very high difficulty due to a low number of very specialized tasks (dealing with classes with functions related to ambient temperature control), three topics have a mean difficulty above $0.45$: Data structure calculations (3) (e.g., computing Pearson correlations between two lists), Text encryption/decryption (14) (e.g., implementing Caesar cipher), and Strings replacements (16) (e.g., replacing a character by another in a string). Compared to HumanEval+, there are more low discriminant topics (discriminant <1<1 or even <0.75<0.75), which highlights that ClassEval contains more tasks where CodeLLMs tend to obtain similar results. As we mentioned in Section 5.2, there were fewer discrepancies between CodeLLMs in terms of the expected probability of answer in ClassEval than HumanEval+. Within these topics, we can even see some of them for which less capable CodeLLMs (Code Gemma and Code Llama) achieve a higher mean accuracy than more capable ones: it is the case for String processing (4) (e.g., removing stop-words from a text with NLTK) or Data structure items additions (15). In the case of Data structure items additions (15) Code Llama and Code Gemma are, on average, better than QwenCoder or GPT3.5. Analysing the tasks within the topic, we observe that more capable models often infer additional assumptions about the task, leading to unnecessarily complex generated code. As an example, we give the generated codes for one such task of GPT3.5 and Code Llama along with the original prompt in Listing 7.
@@
Listing 7: Example of discrepancy on a task from Topic 15. We give first one of our prompt for the task, then generated code by GPT3.5, and CodeLlama.
As one can see, GPT3.5 assumes there is a need to check for some conflicts, possibly due to the ‘course‘ parameter containing information such as ‘start_time‘ and the function ‘self.check_course_conflict‘ being available as a class function from the prompt. However, this is not requested at all in the prompt (and in the task), which is quite straightforward. On the contrary, Code Llama manages the task by following exactly the instructions mentioned in the prompt. We observe a similar phenomenon for other generated code on the same task.
Findings 2: Analysing topics can reveal variations in how CodeLLMs handle different types of tasks. Some topics are notably harder for CodeLLMs in both benchmarks (e.g., Sequences and Nested parentheses/brackets in HumanEval+ and Data structure calculations in ClassEval), while some are trivial (e.g., File Reading in ClassEval). Discriminant information gives us clues on some topics of interest, highlighting the tasks’ topics where CodeLLMs can be more or less easily separated depending on their capabilities.